OpenStudy (anonymous):

A farmer wants to fence an area of 750,000m^2 in a rectangular field and divide it in half with a fence parallel to one of the sides of the rectangle. How can this be done so as to minimize the cost of the fence?

OpenStudy (anonymous):

this on a test go away

OpenStudy (anonymous):

??? It's homework

OpenStudy (rational):

show ur attempt please

OpenStudy (anonymous):

A=area =(L)(W) 750 000=(L)(W) W=750 000/L P=perimeter=3L(2W) P=3L+2(750 000/L) Am I doing this correctly?

OpenStudy (anonymous):

@rational

OpenStudy (rational):

can you sketch the rectangular field you have in mind while setting up those equations

OpenStudy (anonymous):



OpenStudy (anonymous):

Oh I think I got it. So I differentiated the perimeter equation and then set it to 0 and got 0=3L^2-1 500 000 Then after I isolate the L and got sqrt500 000. It then been simplified as 500sqrt2=L

OpenStudy (rational):

looks good to me

OpenStudy (anonymous):

Okay thanks for checking! :D